What is the probability of getting two heads?
If you are talking about coins, 25%
If you are talking about genetics, then it is less than 1%.
What is the meaning of 15 percent precipitation chance on tomorrow?
That out of 100%, there is a 15% chance of rain (if it helps think of it as 15 out of 100 points). So very unlikely it will rain
Three coins are tossed simultaneously the probability of getting atmost one head?
1/2 or 50%
(1/8 prob of 0H + 3/8 prob of 1H)
What is the probability that the first die will have the same number as the second die?
On six sided dice: 1 out of 6. If you make a chart:
- 1 - 2 - 3 - 4 - 5 - 6
1 11 12 13 14 15 16
2 21 22 23 24 25 26
3 31 32 33 34 35 36
4 41 42 43 44 45 46
5 51 52 53 54 55 56
6 61 62 63 64 65 66
See that 6 out of the 36 possibilities (or 1 out of 6) have doubles, meaning that the second die is the same as the first die.
What is the probability of getting a 7 after rolling a single die numbered 1to 6?
What do you think? Its a tiny bit very obvious.
Why can't you use FOIL to multiply a binomial?
you can. i am in algebra II and that's what i was taught
What is the probability of rolling a five with a pair of dice?
There are 36 possible outcomes when we roll a pair of dice. Rolling a five can be done the following ways: 3-2, 4-1, 2-3 and 1-4. That's four ways to roll a five with one roll of the dice. The odds of rolling a five are 4 in 36, or 1 in 9. That makes the probability of rolling a five equal to 0.11111....
We know that the probability of something is a ratio of the number of times it can be done to the total number of things that are possible. Probabilities range from zero to one, inclusive. If something has a probability of zero, it cannot happen. Like rolling a 13. A 13 cannot be rolled with one roll of a pair of standard dice. Thus the probability of rolling a 13 is zero, as it cannot happen.
If something has a probability of one, it must happen. Like the probability of rolling an odd or an even number with one roll of a pair of dice. Since any number we roll must be either odd or even, then the probability of rolling an odd or an even number with one roll of a pair of dice is one. An odd or an even number must turn up, and that is why the probability of rolling an odd or an even number with one roll of the dice is one. Lastly, we've seen that the probability of rolling a given number with one roll of the dice is the number of times that number might appear divided by the total number of possibilities that might appear.
How do you find binomial probabilities using a binomial table?
Read the instructions that accompany the table: they do not all have exactly the same layout.
How many different ways can any two letters of the alphabet be arranged?
Since any two letters of the alphabet do not duplicate each other, then the number of arrangements of any two letters is the same as the number of permutations of 26 things taken 2 at a time. That is 26! / (26 - 2)!, or 26! / 24!, or (26) (25), or 650.
What does it mean for a probability to be fair?
A probability is fair if there is no bias in any of the possible outcomes. Said another way, all of the possible outcomes in a fair distribution have an equal probability.
Yes, as you keep drawing more and more samples and the number of samples become sufficiently large. This is known as the Central Limit Theorem.
What are the odds of being mauled by a tiger?
It likely depends on the tiger, and perhaps how hungry it is.
Is it possible for 2 homozygous recessive individuals to have a homozygous dominant child?
No, homozygous recessive indivuals MUST have two of the reccessive gene (bb), the possible combinations are bb,bb,bb,bb thus only a recessive trait is possible
What is the probability of drawing 4 aces in four draws?
4/52*3/51*2/50*1/49=24/6,497,400
It can be reduced.
1/270,725
Notice the numerator is the number of the desirable choices.
The denominator is the total number available.
Picking the first king: 4/52
Picking the second king: 3/51
Picking the queen of hearts: 1/50
Probability of the whole 3-step maneuver: (4/52) x (3/51) x (1/50) = 12 / 132,600 = 0.00905% (rounded)
The answer to this is 1 minus the probability that they will have 3 or fewer children. This would happen only if they had a boy as the first, second or third child. The probability they have a boy as first child is 0.5
The probability they have a boy as second is 0.25
The probability they have a boy as third is 0.125
Thus the total probability is 0.875
And so the probability they will have more than three children is 1-0.875 or 0.125
The answer depends on what the underlying probability distribution function is.
Most statistical textbooks will contain tables for the binomial, Poisson, Standard Normal, Student's t, Fisher's F, Chi-squared, Pearson's r tests. Then there are lots of non-parametric tests (Spearman's Rank, Mann Whitney U, Kolmogorov-Smirnoff for example) for which there are other tables but these are less widely published. And lastly, there are tests for even more esoteric distributions.
The events in an experiment are mutually exclusive if only one can occur at a time true or false?
provide one business-related example each, with explanation, for mutually exclusive and independent events
What is the probability of rolling 3 3s?
Assuming a standard 6-sided die with numbers 1-6, the probability is:
(1/6) * (1/6) * (1/6) = 1/216 = 0.004629 = 0.4629%
If you are rolling more than three dice, the odds improve dramatically.
It is normal for the ear canal to have some yellowish-brown earwax
What is the probability of getting a black face card from a well shuffled pack of card?
There are 6 black face cards in a standard deck of 52 cards, so the probability of drawing a black face card in one try is 6/52 = 0.115